OpenStudy (anonymous):

I NEED HELP FAST!!! If (6^0)^x = 1, what are the possible values of x? Explain your answer.

OpenStudy (tkhunny):

You tell me. 1? 4? \(\sqrt{2}\)? 0? 1E451?

OpenStudy (anonymous):

i dont know

OpenStudy (tkhunny):

Hogwash. What do you think? \(6^0\) = What?

OpenStudy (anonymous):

1

OpenStudy (tkhunny):

See, you do know! What's \(1^{0}\)? What's \(1^{2}\)? What's \(1^{4}\)? What's \(1^{\sqrt{2}}\)? What's \(1^{5\;million}\)?

OpenStudy (anonymous):

1^0 = 1 1^2 = 2 1^4 = 4 1^ √ 2 = √ 2 1^5 million = 5 million

imqwerty (imqwerty):

:) no this is not correct we will use this rule in our problem->\(\large (m^a)^b=m^{a \times b}\) try using it in this- \(\large (6^0)^x\)

OpenStudy (tkhunny):

Why is 1^2 = 2? I'm pretty sure 1^2 = 1. Try again.

OpenStudy (anonymous):

oh right

OpenStudy (anonymous):

but can we focus on the quetion

OpenStudy (tkhunny):

We are focused on the question. Why is 1^4 = 4? Shouldn't it be 1^4 = 1? We should be seeing a pattern.

OpenStudy (anonymous):

no this isn't helping with the question are just making me answer additional things

OpenStudy (tkhunny):

No, we are looking at material that should lead you to the answer - IF you are paying attention. \(1^{\sqrt{2}} = 1\) Has the result, so far, been anything but "1"?

OpenStudy (anonymous):

no

OpenStudy (tkhunny):

So, if we ALWAYS get 1 - always - what might that mean?